Preferred stock outstanding, 1% 200 800 Net income Average number of shares of common stock outstanding 300The annual report of Sweet Cars, Inc., for the year ended December 31, 2018, included the following items (in millions): (Cli
Answer #1

1)

Preferred dividend = $200*1% = 2

EPS = (Net income-Preferred dividends)/Average number of common shares outstanding

EPS = (800-2)/300

EPS = 798/300

EPS = $2.66

P/E Ratio = market price per share/Earnings per share

P/E Ratio = $35.91/2.66

P/E Ratio = 13.50 times

2. The stock market says that $1 of Fine cars net income is worth = $13.50

Price earnings ratio of 13.50 indicates that to earn $1, Investment needed is $13.50

Hope this helped ! Let me know in case of any queries.